Check whether R is a group under binary operation

a*b=a+b-ab


1
Expert's answer
2021-10-12T10:21:41-0400

(R,*) is a group, if:

1) "\forall a,b\in R: a*b\in R"

2) "\forall a,b\in R: (a*b)*c=a*(b*c)"

3) "\exist e\isin R" "\forall a\in R : a*e=e*a=a"

4) "\forall a\isin R" "\exists a^{-1}\isin R: a*a^{-1}=a^{-1}*a=e"


1) "\forall a,b\in R: (a+b\isin R, ab\in R) \to a*b \in R"

2) "\forall a,b\in R: (a*b)*c = (a+b-ab)*c = a+b-ab+c - (a+b-ab)c ="

"= a+b-ab + c -ac-bc+abc = a-ab-ac+abc+(b+c-bc) =(a+(b+c-bc)) - a(b+c-bc) = a*(b*c)"

3) "a*e = a+e-ae=e+a-ea=e*a"

a * e = a

a + e - ae = a

e(1 - a) = 0

e = 0

"\exist (e = 0)\isin R" "\forall a\in R : a*0=0*a=a"

4) "a*a^{-1} = e"

"a+a^{-1}-aa^{-1} = 0"

"a+a^{-1}(1-a) = 0"

"a^{-1} = -{\frac a {1-a}}"

For a=1 there is no "a^{-1}", so R is not a group under binary operation *.